How do I solve a problem like this: P=795e^0,0306t? with the X that I've got to find at the exponant and not at the base?
Answer by stanbon(48535)

How do I solve a problem like this: P=795e^0,0306t? with the X that I've got to find at the exponent and not at the base?
-------------
P=795e^0,0306t
Divide both sides by 795 to get:
e^(0.0306t) = P/795
-------------
Take the natural log of both sides to get:
0.0306t = ln(P) - ln(795)
-----
t = [ln(P) - ln(795)]/0.0306
=============
Cheers,
Stan H.